# Tilus: A Virtual Machine for Arbitrary Low-Precision GPGPU Computation in LLM Serving

### v2

@@ 6.2. Efficient Low-Precision Support in LLMs
Low-precision kernels in LLMs typically follow two steps before computation: (1) loading weights into on-chip memory (registers or shared memory) from global memory, and (2) casting low-precision weights to high-precision (e.g., float16) followed by de-quantization. Efficient memory loading and casting are thus critical for performance.
@@ 7. Implementation
Tilus comprises five main components: a domain specific language (DSL) in Python, an intermediate representation (IR), optimization passes, a code generator, and a runtime system. The DSL enables developers to write Tilus programs in Python, which are then translated into the VM’s IR for further processing. Optimization passes refine the IR by eliminating redundancies and simplifying arithmetic expressions. The code generator translates the optimized IR into Hidet IR (Ding et al., 2023), a CUDA C-like intermediate representation. Afterwards, we apply the transformations from Section 6 to implement low-precision types with standard precision types while preserving original semantics. The final CUDA C code is generated from Hidet IR and compiled into a hardware binary using the nvcc compiler (NVIDIA Corporation, 2024a). The runtime system manages dynamically loaded binaries and provides the execution environment. The entire system consists of around 20K lines of Python and C++ code. Further details on the compilation pipeline can be found in Appendix C.
@@ 8.1. Experimental Setup
Baselines. We compare our approach, Tilus, against the vendor library cuBLAS (NVIDIA Corporation, 2023), state-of-the-art DL compilers Triton (Tillet et al., 2019) and Ladder (Wang et al., 2024), and hand-crafted kernels QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024). Auto-tuning for Triton (Tillet et al., 2019) and Ladder (Wang et al., 2024) was enabled, while QuantLLM (Xia et al., 2024) used its heuristic policy to select kernel hyper-parameters. For end-to-end evaluations, we integrate our quantized kernels into the state-of-the-art LLM serving framework vLLM (Kwon et al., 2023) and compare them against vLLM (Kwon et al., 2023) and Ladder (Wang et al., 2024) in end to end execution. The specific versions of the tools are: vLLM v0.5.3, Triton v3.1.0, bitblas v0.0.1.dev15 (Ladder), QuantLLM with commit 9802c5a, and Marlin v0.1.1.
@@ 8.2. Performance of Low-Precision Kernels
A single virtual machine program template is implemented to support matrix multiplication with all quantized types, taking tile sizes as tunable hyper-parameters. We denote the performance of this auto-tuned program as Tilus in the evaluation. Figure 9 compares the speedup of Triton (Tillet et al., 2019), Ladder (Wang et al., 2024), QuantLLM (Xia et al., 2024), Marlin (Frantar et al., 2024), and Tilus (ours) against cuBLAS (NVIDIA Corporation, 2023) for various low-precision matrix multiplications: uint8 (u8), float6_e3m2 (f6), uint4 (u4), int4 (i4), uint2 (u2), and uint1 (u1). While each baseline supports a limited set of quantized data types, Tilus consistently achieves speedups across all cases. For small batch sizes, the primary bottleneck is loading weights from global memory to registers for computation on SIMT or Tensor Cores. Triton struggles here due to costly layout conversions after weights are loaded into registers. Although preemptive conversion in global memory could mitigate this, Triton’s programming model lacks explicit layout control, making such optimizations infeasible. Ladder improves upon Triton by modifying data layouts in global memory, avoiding redundant conversions. However, it lacks critical optimizations such as software pipelining (NVIDIA Corporation, 2021; Huang et al., 2023), and its type-level packing limits efficient support for arbitrary bit widths, leading to underutilized memory bandwidth. Expert-crafted kernels from QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024) are optimized for specific quantization schemes but lack flexibility and maintainability. In contrast, Tilus outperforms all baselines with a single parameterized virtual machine template, efficiently supporting a full range of quantization types through a well-abstracted programming model.
@@ 8.3. Arbitrary Data Type Support
Our virtual machine (VM) supports low-precision matrix multiplications of the form matmul(A, B), where operand A can have data types with 32, 16, or 8 bits, and weight B supports a wide range of bit widths, from 32 bits down to 1 bit. Standard data types such as float32, float16, and int8 are supported, along with customized low-precision types with fewer than 8 bits, which include signed integers, unsigned integers, and floating-point formats with arbitrary exponent and mantissa distributions. Leveraging the algebraic layout system (Section 4 and 6.2), our VM enables efficient memory access for low-precision data. Figure 10 illustrates the speedup achieved for the full spectrum of quantized weight data type: uint1 to uint8, int2 to int8, and float3 to float8. Representative exponent-mantissa distribution of floating-point data types such as e4m3, e3m3, e3m2, e2m2, e2m1, and e1m1 are chosen. Each row represents the type kind (e.g., unsigned integer, signed integer or floating data type) while each columns represents the bit width. We take 128 as the group size of sub-channel quantization. Using matrix multiplication dimensions of BS=16, K=8192, and N=57344 the results demonstrate substantial speedups. These findings validate that our VM can effectively support arbitrary low-precision types with high efficiency, making it a robust solution for low-precision computations in modern GPUs. We want to highlight that all the kernels are from the same program template, by parameterizing tile sizes, thus the programming efforts are limited. There are around 200 configurations per operator, and it takes around one minute to compile. We used float16 as the activation data type in the experiment and we also support bfloat16 and int8.
@@ 8.4. End-to-End Performance
We evaluated the end-to-end performance of representative LLMs: Gemma-2-9B (Team et al., 2024), QWen-2.5-32B (Yang et al., 2024), and Llama-3.3-70B (Grattafiori et al., 2024), across both prefill and decode stages. The prefill stage processes all prompt tokens at once, generating the kv-cache for subsequent token generation during the decode stage, which iteratively generates one token at a time. Prefill latency determines the time-to-first-token (TTFT), while decode latency impacts the speed of subsequent token generation. Both stages are critical for optimizing user experience and system utilization. Contiguous batching (Yu et al., 2022; Kwon et al., 2023) was used to batch multiple decode requests efficiently. Figure 11 shows the latency of both stages across these models. Our method consistently outperforms Ladder (Wang et al., 2024), particularly in the decode stage with more than one token (middle column of Figure 11). Analysis of Ladder’s generated kernels revealed suboptimal use of CUDA Cores for 1–15 tokens and Tensor Cores for 16 or more tokens, as key optimizations like software pipelining (Huang et al., 2023) and k-dimension parallelization (Osama et al., 2023) were not implemented, leading to poor performance. For the prefill stage, we decode quantized weights to float16 and perform computations using standard f16xf16 matrix multiplication kernels, as computation becomes the bottleneck at this stage. Our efficient handling of quantized weight layouts ensures minimal overhead for decoding, contributing to the superior performance observed.
@@ 8.5.1. Speedup over Different Hardware
We evaluate the end-to-end performance of the QWen2.5-30B model on NVIDIA A100, L40S, and H100 GPUs, which correspond to the Ampere, Ada Lovelace, and Hopper architectures, respectively. Figure 12 presents a performance comparison of vLLM (Kwon et al., 2023) (float16), Ladder (Wang et al., 2024) (uint4), and Tilus (uint4, ours) across the decode and prefill stages. On the Hopper architecture (H100), Ladder fails to generate valid kernels, leading to a CUDA error (’an illegal instruction was encountered’), which we denote as ERR in the figure. On the L40S GPU, vLLM (Kwon et al., 2023) exceeds the available 48 GiB DRAM capacity, leading to out-of-memory (OOM) errors. In all other configurations, Tilus consistently outperforms Ladder across all GPUs and both processing stages, highlighting its robust performance and adaptability across architectures.
@@ 8.5.2. Speedup over Different Batch Sizes
We analyze the relationship between speedup and batch size by benchmarking matrix multiplication performance under different batch sizes. For the decode stage, we evaluate batch sizes of 1, 4, 8, and 16, while for the prefill stage, we use batch sizes of 4096, 8192, and 12,288. Experiments are conducted on Llama-3.3-70B-Instruct (Grattafiori et al., 2024) model with quantized data types float6_e3m2 (f6) and uint4 (u4), using $k=8192$ and $n=57344$. As shown in Figure 13, Tilus consistently outperforms baseline approaches across all batch sizes that are used in both decode and prefill stages of LLM serving. To further understand these performance differences, we profile kernels from cuBLAS, Ladder, and Tilus, with detailed analysis provided in Appendix D.
@@ 9. Related Work
Many deep learning compilers adopt loop-oriented scheduling (Chen et al., 2018; Ragan-Kelley et al., 2013) and build auto-tuning frameworks on top of it (Adams et al., 2019; Vaswani et al., 2017; Vasilache et al., 2018; Ahn et al., 2020; Xing et al., 2022; Zheng et al., 2022b, 2020a, a, 2020b; Shao et al., 2024; Feng et al., 2023; Wang et al., 2024). In contrast, Tilus employs a procedure-oriented approach that better models GPU hardware, improving programmability and flexibility. Beyond loop-oriented scheduling, tensor programs are often optimized using vendor libraries (e.g., cuBLAS (NVIDIA Corporation, 2023)), predefined templates for efficient matrix multiplication (NVIDIA Corporation, 2021), hardware-aware tiling strategies (Zhu et al., 2022), and domain-specific compilers for linear algebra (Sabne, 2020). While these methods prioritize performance, they lack extensibility for arbitrary low-precision data types. Other research focuses on optimizing irregular or ragged tensor programs (Tang et al., 2022; Fegade et al., 2022), operator fusion (Zheng et al., 2022c; Zhao et al., 2022), dynamic shape handling (Shen et al., 2021; Zhu et al., 2021; Fegade et al., 2021; Zheng et al., 2022b), and scheduling independent operators (Ma et al., 2020; Ding et al., 2021; Jia et al., 2019). Microscaling data types (Rouhani et al., 2023) can be thought as a more fine-grained quantization thus we could also support it. These techniques are complementary to our focus on efficient low-precision computation. Triton (Tillet et al., 2019) introduces a tile-based programming model. However, it lacks explicit support for low-precision data types and does not expose the GPU memory hierarchy, limiting optimization opportunities. Similarly, Hidet (Ding et al., 2023), which serves as our backend, does not provide built-in support for low-precision types. Graphene (Hagedorn et al., 2023) presents an intermediate representation (IR) with a layout representation. Unlike Graphene’s focus on strides and computation, our algebraic layout system emphasizes hierarchical organization. In fact, we can express Graphene’s layout representation as a primitive component within our system.
@@ 10. Conclusion
We introduced Tilus, a GPGPU virtual machine designed for efficient low-precision LLM serving, addressing key limitations of existing solutions. Tilus features an algebraic layout system for tensor distribution within thread block registers, a thread-block-level programming model with fine-grained memory management, and broad support for sub-byte data types, enabling arbitrary precision from 1 to 8 bits. Experimental results show substantial performance gains over state-of-the-art frameworks like Triton and Ladder, demonstrating the flexibility and scalability of our approach. This work establishes a foundation for efficient and extensible LLM inference, paving the way for further optimizations in emerging hardware, advanced quantization techniques, and diverse low-precision formats.

### v3

@@ 7.2. Efficient Low-Precision Support in LLMs
[…]
@@ 8. Implementation
Tilus comprises five main components: a domain specific language (DSL) in Python, an intermediate representation (IR), optimization passes, a code generator, and a runtime system. The DSL enables developers to write Tilus programs in Python, which are then translated into the VM’s IR for further processing. Optimization passes refine the IR by eliminating redundancies and simplifying arithmetic expressions. The code generator translates the optimized IR into Hidet IR (Ding et al., 2023), a CUDA C-like intermediate representation. Subsequently, we apply the transformations from Section 7 to implement low-precision operations using standard precision types while preserving original semantics. The final CUDA C code is generated from Hidet IR and compiled into a hardware binary using the nvcc compiler (NVIDIA Corporation, 2024a). The runtime system manages dynamically loaded binaries and provides the execution environment. The entire system consists of approximately 35K lines of Python and C++ code.
8.1. Program Compilation and Runtime
Tilus provides a user-friendly interface that allows programmers to write Tilus programs directly in Python, simplifying the integration of produced kernels with the rich deep learning ecosystem. Given a program, we take several steps to compile it to GPU executable code.
Step 1: Global and Shared Memory Planning Each GPU kernel can use a shared memory space of a size determined at launch-time. To simplify GPU programming, we allow the users to allocate shared memory multiple times in the program on demand. Thus, we need a shared memory planner to calculate the size of shared memory the Tilus program needs, and map the shared tensor to one region of the kernel’s shared memory space. Similar to shared memory planning, we also require a global memory planner to manage the allocation of global memory shared by all thread blocks. We request Tilus’s runtime system to allocate a workspace in global memory, enabling the kernel to use this workspace via the AllocateGlobal instruction during its execution.
Step 2: Code Emitting for Each Instruction. We emit low-level GPU code for each Tilus instruction sequentially. In our implementation, we use the Hidet IR (Ding et al., 2023) to represent the low-level GPU code. During this process, we perform instruction selection to choose the most efficient low-level instructions where feasible. For example, we use lds PTX instruction (NVIDIA Corporation, 2024b) to load the data from shared memory to register. However, a more efficient PTX instruction ldmatrix could also be used if the layout of the loaded register tensor is compatible with the layout spatial(8, 4).repeat(1, 4), which means it can be divided properly by the latter. Besides, we also perform automatic vectorization for memory loading and storing instructions. For example, we use vectorized instructions, such as cp.async.v4, lds128, and ldg128, to maximize memory access efficiency.
Step 3: Lowering Low-Precision Data Types. After we emit thread-block-level instructions to the low-level IR, we will apply the passes that implement the the rules discussed in Section 8 to transform all low-precision operations in the low-level IR to corresponding operations on hardware-friendly types. In most cases, only the vectorized type casting from low-precision type to standard type (e.g., float16) will be applied since the memory loading of low-precision data will be replaced by standard types thanks to our layout formalization and register tensor reinterpretation. Subsequently, we generate CUDA code (for NVIDIA GPUs) from the low-level IR and then use the nvcc compiler to produce the hardware binary that can be dynamically loaded.
Step 4: Loading by Runtime System. The compiled binary is dynamically loaded by the Tilus’s runtime system. The runtime system also maintains internal states to serve the kernel execution: 1) a workspace memory allocated on-demand, which compiled kernels can request via AllocateGlobal instruction; 2) an execution context that stores the CUDA stream on which the kernel are launched; and 3) the kernels cached in memory to avoid unnecessary recompilation.
@@ 9.1. Experimental Setup
Baselines. We compare our approach, Tilus, against the vendor library cuBLAS (NVIDIA Corporation, 2023), state-of-the-art DL compilers Triton (Tillet et al., 2019) and Ladder (Wang et al., 2024), and hand-crafted kernels QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024). Auto-tuning for Triton (Tillet et al., 2019) and Ladder (Wang et al., 2024) was enabled, while QuantLLM (Xia et al., 2024) used its heuristic policy to select kernel hyperparameters. For end-to-end evaluations, we integrate our quantized kernels into the state-of-the-art LLM serving framework vLLM (Kwon et al., 2023) and compare them against vLLM (Kwon et al., 2023) and Ladder (Wang et al., 2024) in end to end execution. The specific versions of the tools are: vLLM v0.5.3, Triton v3.1.0, bitblas v0.0.1.dev15 (Ladder), QuantLLM with commit 9802c5a, and Marlin v0.1.1.
@@ 9.2. Performance of Low-Precision Kernels
A single virtual machine program template is implemented to support matrix multiplication with all quantized types, taking tile sizes as tunable hyperparameters. We denote the performance of this auto-tuned program as Tilus in the evaluation. Figure 10 compares the speedup of Triton (Tillet et al., 2019), Ladder (Wang et al., 2024), QuantLLM (Xia et al., 2024), Marlin (Frantar et al., 2024), and Tilus (ours) against cuBLAS (NVIDIA Corporation, 2023) for various low-precision matrix multiplications: uint8 (u8), float6_e3m2 (f6), uint4 (u4), int4 (i4), uint2 (u2), and uint1 (u1). While each baseline supports a limited set of quantized data types, Tilus consistently achieves speedups across all cases. For small batch sizes, the primary bottleneck is loading weights from global memory to registers for computation on SIMT or Tensor Cores. Triton struggles here due to costly layout conversions after weights are loaded into registers. Although changing layout in global memory could mitigate this, Triton’s programming model lacks explicit layout control, making such optimizations infeasible. Ladder improves upon Triton by modifying data layouts in global memory, avoiding redundant conversions. However, it lacks critical optimizations such as software pipelining (NVIDIA Corporation, 2021; Huang et al., 2023), and its type-level packing limits efficient support for arbitrary bit widths, leading to underutilized memory bandwidth. Expert-crafted kernels from QuantLLM (Xia et al., 2024) and Marlin (Frantar et al., 2024) are optimized for specific quantization schemes but lack flexibility and maintainability. In contrast, Tilus consistently outperforms all baselines using a single parameterized Tilus program template, which efficiently supports a full range of quantization types through a well-abstracted programming model.
@@ 9.3. Arbitrary Data Type Support
[…]
Tilus supports low-precision matrix multiplications of the form matmul(A, B), where operand A can have data types with 32, 16, or 8 bits, and weight B supports a wide range of bit widths, from 32 bits down to 1 bit. Standard data types such as float32, float16, and int8 are supported, along with customized low-precision types with fewer than 8 bits, which include signed integers, unsigned integers, and floating-point formats with arbitrary exponent and mantissa distributions. Leveraging the algebraic layout system (Section 4 and 7.2), Tilus enables efficient memory access for low-precision data. Figure 11 illustrates the speedup achieved for the full spectrum of quantized weight data types: uint1 to uint8, int2 to int8, and float3 to float8. Representative exponent-mantissa distribution of floating-point data types such as e4m3, e3m3, e3m2, e2m2, e2m1, and e1m1 are chosen. Each row represents the type kind (e.g., unsigned integer, signed integer or floating data type) while each column represents the bit width. Using matrix multiplication dimensions of BS=16, K=8192, and N=57344 the results demonstrate substantial speedups. These findings validate Tilus’s effectiveness in supporting arbitrary low-precision types with high efficiency, making it a robust solution for low-precision computations in modern GPUs. Notably, all kernels are generated from the same program template by parameterizing tile sizes, which limits the required programming effort. There are around 200 configurations per operator, and it takes around one minute to compile. We used float16 as the activation data type in the experiment and we also support bfloat16 and int8.
@@ 9.4. End-to-End Performance
[…]
We evaluated the end-to-end performance of representative LLMs: Gemma-2-9B (Team et al., 2024), QWen-2.5-32B (Yang et al., 2024), and Llama-3.3-70B (Grattafiori et al., 2024), across both prefill and decode stages. The prefill stage processes all prompt tokens at once, generating the kv-cache for subsequent token generation. The decode stage then iteratively generates one token at a time. Prefill latency determines the time-to-first-token (TTFT), while decode latency impacts the speed of subsequent token generation. Both stages are critical for optimizing user experience and system utilization. Contiguous batching (Yu et al., 2022; Kwon et al., 2023) was used to efficiently batch multiple decode requests. Figure 12 shows the latency of both stages across these models. Our method consistently outperforms Ladder (Wang et al., 2024), particularly in the decode stage for batch sizes greater than one (middle column of Figure 12). Analysis of Ladder’s generated kernels revealed suboptimal use of CUDA Cores for 1–15 tokens and Tensor Cores for 16 or more tokens, as key optimizations like software pipelining (Huang et al., 2023) and k-dimension parallelization (Osama et al., 2023) were not implemented, leading to poor performance. For the prefill stage, quantized weights are decoded to float16, and computations are performed using standard f16xf16 matrix multiplication kernels, as computation becomes the bottleneck at this stage. Our efficient handling of quantized weight layouts ensures minimal overhead for decoding, contributing to the superior performance observed.
@@ 9.5.1. Speedup over Different Hardware
[…]
We evaluate the end-to-end performance of the QWen2.5-30B model on NVIDIA A100, L40S, and H100 GPUs, which correspond to the Ampere, Ada Lovelace, and Hopper architectures, respectively. Figure 13 presents a performance comparison of vLLM (Kwon et al., 2023) (float16), Ladder (Wang et al., 2024) (uint4), and Tilus (uint4, ours) across the decode and prefill stages. On the Hopper architecture (H100), Ladder is unable to generate valid kernels, leading to a CUDA error (’an illegal instruction was encountered’), which we denote as ERR in the figure. On the L40S GPU, vLLM (Kwon et al., 2023) exceeds the available 48 GiB DRAM capacity, leading to out-of-memory (OOM) errors. In all other configurations, Tilus consistently outperforms Ladder across all GPUs and both processing stages, highlighting its robust performance and adaptability across architectures.
@@ 9.5.2. Speedup over Different Batch Sizes
[…]
We analyze the relationship between speedup and batch size by benchmarking matrix multiplication performance under different batch sizes. For the decode stage, we evaluate batch sizes of 1, 4, 8, and 16, while for the prefill stage, we use batch sizes of 4096, 8192, and 12,288. The batch size corresponds to the number of tokens processed in one step. In the prefill stage, it equals the sum of sequence length of all requests, while in decode stage, it equals to the number of requests since each request only generates one token each time. Experiments are conducted on Llama-3.3-70B-Instruct (Grattafiori et al., 2024) model with quantized data types float6_e3m2 (f6) and uint4 (u4), using $k=8192$ and $n=57344$. As shown in Figure 14, Tilus consistently outperforms baselines across all batch sizes that are used in both decode and prefill stages of LLM serving.
@@ 10. Related Work
Many deep learning compilers adopt loop-oriented scheduling (Chen et al., 2018; Ragan-Kelley et al., 2013) and build auto-tuning frameworks on top of it (Adams et al., 2019; Vaswani et al., 2017b; Vasilache et al., 2018; Ahn et al., 2020; Xing et al., 2022; Zheng et al., 2022b, 2020a, a, 2020b; Shao et al., 2024; Feng et al., 2023; Wang et al., 2024). In contrast, Tilus employs a procedure-oriented approach that better models GPU hardware, improving programmability and flexibility. Beyond loop-oriented scheduling, tensor programs are often optimized using vendor libraries (e.g., cuBLAS (NVIDIA Corporation, 2023)), predefined templates for efficient matrix multiplication (NVIDIA Corporation, 2021), hardware-aware tiling strategies (Zhu et al., 2022), and domain-specific compilers for linear algebra (Sabne, 2020). While these methods prioritize performance, they lack extensibility for arbitrary low-precision data types. Other research focuses on optimizing irregular or ragged tensor programs (Tang et al., 2022; Fegade et al., 2022), operator fusion (Zheng et al., 2022c; Zhao et al., 2022), dynamic shape handling (Shen et al., 2021; Zhu et al., 2021; Fegade et al., 2021; Zheng et al., 2022b), and scheduling independent operators (Ma et al., 2020; Ding et al., 2021; Jia et al., 2019). Prior works (Zeng et al., 2025; Ma et al., 2025; Feng et al., 2021; Park et al., 2024) endeavor to support quantization with arbitrary bit-width by independently processing different bits. However, these methods are primarily limited to integer quantization and often exhibit suboptimal execution performance. PartIR (Alabed et al., 2025) also introduces a layout system for tiles; however, its abstraction level is higher than Tilus’s, and it does not detail the assignment of tile elements to threads. ExTensor (Hegde et al., 2019) and FuseMax (Nayak et al., 2024) leverage einsum to define computations, implicitly determining memory access patterns. Nevertheless, they do not specify how tile elements are distributed among GPU threads. Tilus can also accommodate codebook quantization, such as LCQ (Cai et al., 2024), to facilitate low-precision computation through the addition of a new lookup instruction. Microscaling data types (Rouhani et al., 2023) can be thought as a more fine-grained quantization thus we could also support it. These techniques are complementary to our focus on efficient low-precision computation. Triton (Tillet et al., 2019) introduces a tile-based programming model. However, it lacks explicit support for low-precision data types and does not expose the GPU memory hierarchy, limiting optimization opportunities. Similarly, Hidet (Ding et al., 2023), which serves as our backend, does not provide built-in support for low-precision types. Graphene (Hagedorn et al., 2023) presents an intermediate representation (IR) with a layout representation. Unlike Graphene’s focus on strides and computation, our algebraic layout system emphasizes hierarchical organization. In fact, we can express Graphene’s layout representation as a primitive component within our system.
@@ 11. Conclusion
We introduced Tilus, a tile-level GPGPU programming language designed to expose shared memory and registers to developers, which enables the creation of efficient low-precision kernels for LLM serving. Tilus features an algebraic layout system for managing tensor distribution across threads, a thread-block-level programming model with fine-grained memory management, and comprehensive support for sub-byte data types, enabling arbitrary precision from 1 to 8 bits. Our experimental results demonstrate substantial performance gains over state-of-the-art approaches, showcasing the flexibility and scalability of our method. This work establishes a foundation for efficient and extensible LLM inference, paving the way for future optimizations in emerging hardware, advanced quantization techniques, and diverse low-precision formats.
Acknowledgement
We would like to thank the members of the EcoSystem research laboratory at the University of Toronto for their feedback on the early manuscript. We also thank the anonymous ASPLOS reviewers and our shepherd, Jacques Pienaar, for their valuable feedback and suggestions, as well as the artifact evaluation reviewers for reproducing our experiments. The authors affiliated with the University of Toronto were supported by the Canada Foundation for Innovation JELF grant, the NSERC Discovery grant, the Google Scholar Research Award, the VMware Early Career Faculty Grant, and the Ontario Early Researcher Award.